OpenStudy (anonymous):

complete the factoring of (20t)(20t)-15t=5t(?)

OpenStudy (campbell_st):

it can be written as \[(5t \times 4)(20t) - (5t \times 3) = 5t(?)... or ... 5t(4 \times 20t) - 5t(3) = 5t (?)\] hopefully this helps in identifying what should be inside the brackets

OpenStudy (mertsj):

Or perhaps \[20t(20t-1)=0\]

OpenStudy (campbell_st):

what happened to (?)

OpenStudy (mertsj):

Oh. I see your point. I didn't realize that was part of the problem. Thought the asker added it.

OpenStudy (mertsj):

So perhaps 5t(80t-3)

OpenStudy (campbell_st):

Me thinks so...

OpenStudy (mertsj):

I think you and I are the only ones who care.